you have 30 coins how many nickles and dimes does it take to make $2.10
Found 2 solutions by ikleyn, Boreal:
Answer by ikleyn(52803)

.

5N + 10*(30-N) = 210 cents


5N - 10N = 210 - 10*30

   -5N = -90

     N = 18.


ANSWER.  18 nickels and 12 dimes.



Answer by Boreal(15235)

N+D=30 count equation
.05N+.10D=2.10 money equation
N=30-D substitute
.05(30-D)+.10(D)=2.10
1.5-.05D+.10D=2.10
0.5D=0.60
D=12
N=18
$1.20+$0.90=$2.10